OpenStudy (anonymous):

8. A ball is dropped from a height of 160 cm. Each time it drops it rebounds to 80% of its previous height. b) What is the total distance travelled at the time of the 10th bounce?

OpenStudy (unklerhaukus):

\[160 + \sum_{n=1}^{10} 160 \times 2 \times 0.8^n\]

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

UnkleRahukas It is an infinite series

OpenStudy (unklerhaukus):

i thought there were only 10 bounces?

OpenStudy (anonymous):

the answer is 1268.2m just so to let you guys know Im looking for a sum

OpenStudy (anonymous):

Sorry i calculated the total distance that it would travel it is definitely not 1268.2m

OpenStudy (anonymous):

what did you get?

OpenStudy (unklerhaukus):

even if it bounced 100% of original bounce you will not get a total distance greater than 21x1.6m =34m

OpenStudy (anonymous):

i think they meant sum find the sum of all the distances

OpenStudy (unklerhaukus):

well i dont know if i have done this correctly but i have ~\[ 160+160 \times 2 \times 4.57\]~\[16.22meters\]